OpenStudy (anonymous):

How do I integrate (sin^2x-cos^2x)/sinx?

OpenStudy (anonymous):

fair...lol

OpenStudy (anonymous):

sinx - cosx*cotx

OpenStudy (dumbcow):

you could also get it in terms of just sin cos^2 = 1-sin^2

OpenStudy (anonymous):

sinx - (1-sin^2x)/sinx = 2sinx -1/sinx

OpenStudy (anonymous):

I think you can do it from here

OpenStudy (anonymous):

i hav triied it uhh..uh

OpenStudy (anonymous):

Ruchi what did you just paste?

OpenStudy (anonymous):

i need an indefinite integral... how did you even get to a numerical solution on an indefinite integral?

OpenStudy (anonymous):

you have to integrate it

OpenStudy (anonymous):

Ignore that, do what dumbcow told you

OpenStudy (anonymous):

yeahhhhhhhhhhhpii

OpenStudy (anonymous):

yeah im doing the work on that now... ill get back if i need more help from there

OpenStudy (anonymous):

zacha.. m i gona get it ahh?

OpenStudy (anonymous):

ruchi you told me nothing...

OpenStudy (anonymous):

(sinx)^2+(cosx)^2=1 or (cosx)^2=1-(sinx)^2 So: 1-(sinx)^2-(sinx)^2=sin(x) 1-2(sinx)^2=sin(x) 0=2(sinx)^2+sin(x)-1 sinx=[-1+sqrt(1+8)]/4 or [-1-sqrt(1+8)]/4 sinx=[-1+3]/4 or [-1-3]/4 sinx=2/4 or -4/4 sinx=1/2 or -1 If sinx=0.5 x=pi/6+2piK or x=5pi/6+2piK If sinx=-1 x=-pi/2+2piK All the possible solutions are actually x=pi/6+(2/3)piK And -pi<x<=pi The solutions are: x= -pi/2, pi/6, 5pi/6

OpenStudy (anonymous):

that's solving for x if it were an equation... i'm looking for the INDEFINITE INTEGRAL.

OpenStudy (anonymous):

so i simplified and got integral(3sinx-cscx) which= ln|cscx+cotx|-2cosx+C

OpenStudy (anonymous):

sorry that should be a 2 in front of the sin

OpenStudy (anonymous):

integrate (sin^2x-cos^2x)/2sinx -( cos^2x - sin^2x) = -(1 - 2sin^2x) = 2sin^2x -1 ->(sin^2x-cos^2x)/2sinx = sinx - 1/2sinx Int (sinx - 1/2sinx) dx = -cosx + ln |cscx + cotx| + c

OpenStudy (anonymous):

OOps, missing 2 at Ln. = -cosx + (1/2) ln |cscx + cotx| + c